A juice company purchases components from a variety of sources through their global network of suppliers. They are looking to determine their order quantity for PET bottles. They have forecast their demand for PET bottles for the next year to be 1206115 cases. Being a standard component, sales are expected to be constant/uniform over the course of the year. They operate 5 days per week so each year is made up of 260 days. The lead time for PET bottles (time from order to receipt is 2 days. They have also determined that the cost to carry a case of PET bottles in inventory is $1.38 per case per year and that order costs are estimated at $78.96 per order.  (Note: when working with inventory problems demand and carrying costs must be stated in the same time frame, annual in this case) The PET bottles purchased for $1.24 per case.  Given that the assumptions for the basic EOQ model apply, calculate the optimal order quantity (in cases) for this item. How many days will each order last?
Answer #1

Given:Annual demand=1206115

Number of working days =260;

Lead time =2 days

holding cost per unit per year=$ 1.38

cost per order=$ 78.96

Optimal order quantity also known as Economic order quantity is given by the formula ,

EOQ=sqrt(2*Annual demand*cost per order/holding cost per unit per year)

=sqrt(2*1206115*78.96/1.38)

=11748.256

=11748 cases per order (rounding up)

Number of orders =Annual demand/EOQ=1206115/11748=102.6656 =103 orders per year

Time between orders =Number of working days /Number of orders

=260/103

=2.524

=2.5 days

Number of days each order will last=time between orders =2.5 days